A student tossed three coins. What is the probability of getting at least two heads? This is a Probabilty of independant events 1/2 * 1/2 = 1/4 or 25%
Found 2 solutions by stanbon, G8TORS:
Answer by stanbon(57255)

A student tossed three coins. What is the probability
of getting at least two heads?
----
P(2<= x <=3) = 3C2(1/2)^3 + 3C3(1/2)^3 = (3+1)/8 = 1/2
====================
Cheers,
Stan H.

Answer by G8TORS(10)

A student tossed three coins. What is the probability of getting at least two heads?
This is a Probabilty of independant events 1/2 * 1/2 = 1/4 or 25%